# Multi-Objective Coverage Bayesian Optimization (MOCOBO)

### v3

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In this work, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “covers" the $T$ objectives. A set of solutions is defined as “covering" if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. To address this problem, we propose Multi-Objective Coverage Bayesian Optimization (MOCOBO), a principled algorithm designed to efficiently find a covering set. We validate our approach through experiments on challenging high-dimensional tasks, including applications in peptide and molecular design, where MOCOBO is shown to find high-performing covering sets of solutions. The results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery. We make code available here: https://github.com/nataliemaus/mocobo.
@@ 1 Introduction
Figure 1: Traditional multi-objective optimization for $T$ objectives might select any point along the Pareto frontier, but in some situations like this any Pareto optimal point performs poorly on at least one objective. In situations where multiple $K<T$ solutions are allowed ($\blacksquare$), we can sometimes optimize all objectives well.
@@
Bayesian optimization (BO; 31, 54, 21) is a general framework for sample-efficient optimiziation of black-box functions. By using a probabilistic surrogate model, such as a Gaussian process, Bayesian optimization balances exploration and exploitation to identify high-performing solutions with a limited number of function evaluations. BO has been successfully applied in a wide range of domains, including hyperparameter tuning [56, 64], A/B testing [37], chemical engineering [26], drug discovery [47], and more.
In multi-objective optimization the typical assumption is that objectives inherently trade off with one another, motivating the search for Pareto-optimal solutions that balance these trade-offs. Multi-objective Bayesian optimization methods focus on identifying diverse (in objective value) sets of solutions along the Pareto front to capture these trade-offs [27, 3, 63, 33, 11, 57, 4]. Indeed, without such trade-offs, optimization would trivialize into simply optimizing all of the objectives.
However, in many real-world scenarios, the trade-offs between objectives can become extreme, with some objectives being nearly mutually exclusive. This makes it virtually impossible to achieve satisfactory performance across all objectives simultaneously. Figure˜1 provides a $1D$ example of this demonstrating how the Pareto-optimal points may fail to achieve acceptable performance on all objectives. Such scenarios motivate an interesting generalization of the multi-objective setting that arises in situations where we can propose a small set of solutions that together optimize the multiple objectives and avoid the implicit constraints imposed by extreme trade-offs.
One compelling example arises in drug design. In this domain, the goal may be to develop antibiotics that treat a broad spectrum of pathogens–here, activity against the individual pathogens serve as the objectives. While it is very true that we seek to optimize the objectives simultaneously in the sense that we prefer to have broad-spectrum antibiotics, it is at the same time perfectly reasonable for one drug in the set to be ineffective against a few of the pathogens if it allows for greater potency against others. This flexibility enables a more targeted and efficient approach to therapeutic design, where our realistic goal is to have a small set of drugs that are together extremely lethal for the set of pathogens, without requiring each individual drug to be.
To formalize this, we consider the “coverage optimization" problem: the task of covering a set of $T$ objectives with $K$ solutions, where coverage means that for each objective $f_{1},\dots,f_{T}$, there exists at least one high-performing solution among the $K$. In the context of drug design, this translates to identifying $K$ drugs such that each of the $T$ pathogens is effectively addressed by at least one drug in the set. This problem setting diverges from traditional multi-objective optimization by prioritizing collective coverage over balancing trade-offs within individual solutions.
To address the coverage problem setting, we introduce and experimentally validate Multi-Objective Coverage Bayesian Optimization (MOCOBO), a new algorithm designed to solve the “coverage optimization" problem within the Bayesian optimization framework. MOCOBO efficiently identifies a small set of $K$ solutions that collectively cover $T$ objectives. Unlike traditional multi-objective Bayesian optimization methods that seek to balance trade-offs across all objectives for every solution, MOCOBO focuses on identifying specialized sets of solutions to meet practical coverage requirements.
Contributions
  1. 1.
    We consider the problem setting of finding a “covering set" of $K<T$ high-performing solutions such that each relevant objective $f_{t}\in\left\{f_{1},\ldots f_{T}\right\}$ is optimized by at least one solution. As far as we are aware, this is the first work in the Bayesian optimization (BO) literature to consider the coverage optimization problem setting.
  2. 2.
@@
    We propose a local BO algorithm, MOCOBO, for this problem setting, that extends to high dimensional objectives, structured input domains, and large numbers of function evaluations.
  3. 3.
@@
2 Background and related work
Bayesian optimization (BO).
BO is a method for sample-efficient black-box optimization. It operates iteratively, where each cycle involves training a probabilistic surrogate model, typically a Gaussian process (GP; 50), on data collected from the black-box objective function. BO then employs the predictive posterior $p(y\mid\mathbf{x},D)$ of the surrogate model to compute an acquisition function that gives a policy for choosing the next candidate(s) to evaluate, balancing exploration and exploitation.
Multi-Objective Bayesian optimization (MOBO).
Recent advances in MOBO have introduced numerous methods aimed at identifying diverse solutions along the Pareto front to balance trade-offs among multiple objectives [27, 3, 63, 33, 11, 57, 4]. These approaches, however, are not well-suited to “coverage optimization”, where the goal is to find a small set of $K<T$ solutions that collectively address all $T$ objectives. Despite this, we compare our proposed method to MORBO, developed by Daulton et al. [11], which is regarded as a state-of-the-art method for high-dimensional MOBO tasks.
Bayesian optimization over structured search spaces.
BO has recently been used for optimizing structured search spaces like molecules or amino acid sequences by employing latent space Bayesian optimization, which utilizes a variational autoencoder (VAE; 32, 52) to convert structured inputs into a continuous latent space for BO [23, 15, 61, 24, 55, 30, 57]. The VAE encoder, $\Phi(\mathbf{x})$, transforms structured inputs $\mathbf{x}$ into continuous latent vectors $\mathbf{z}$, enabling BO to operate in the latent space. Candidate latent vectors are decoded using the VAE decoder, $\Gamma(\mathbf{z})$, to generate structured outputs. Maus et al. [44] extended TuRBO [17] for latent space optimization by jointly training the surrogate model and VAE using variational inference.
@@ 2 Background and related work
Local Bayesian optimization [17], including TuRBO-$M$, is one recent successful approach to high-dimensional BO. TuRBO-$M$ uses $M$ parallel local optimization runs, each maintaining its own dataset $\mathcal{D}_{i}$ and surrogate model. Each local optimizer proposes candidates only within a hyper-rectangular trust region $\mathcal{T}_{i}$. The trust region $\mathcal{T}_{i}$ is a rectangular area of the input space $\mathcal{X}$, centered on the incumbent $\mathbf{x}^{+}_{i}$. Its side length $\ell_{i}$ is bound by $\ell_{min},\ell_{max}]$. If an optimizer improves the incumbent for $\rho_{succ}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{max})$. If not improved in $\rho_{fail}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{min}$. While TuRBO-$M$ is not directly applicable to our problem, we follow methods like MORBO [11] and ROBOT [45] in adapting the use of $M$ coordinated trust regions ($K$ in our case) to a different problem setting.
Prior work on coverage optimization and related problem settings.
To our knowledge, the coverage optimization problem setting considered in this paper has not been previously explored in the Bayesian optimization literature. However, similar problem settings have been studied in the context of gradient-based multi-objective optimization. In particular, Ding et al. [13], Li et al. [39], and Lin et al. [40] have investigated related problems in this domain. Ding et al. [13] demonstrate that the coverage optimization problem generalizes various clustering methods, including k-means clustering, and propose an algorithm that extends k-means++ [1] and Lloyd’s algorithm [43] to efficiently solve the coverage problem using gradient descent. Li et al. [39] introduce Many-objective Multi-solution Transport (MosT), a novel framework that scales multi-objective gradient-based optimization to a large number of objectives. MosT uses an optimal transport to establish a balanced assignment between objectives and solutions, allowing each solution to focus on a specific subset of objectives and ensuring that the collective set of solutions covers all objectives. Lin et al. [40] introduce a novel Tchebycheff set (TCH-Set) scalarization approach for covering multiple conflicting objectives using a small set of collaborative solutions. They further propose a smooth Tchebycheff set (STCH-Set) scalarization method to handle non-smoothness in TCH-Set scalarization. While these three papers address related coverage optimization problems, their proposed methods rely on differentiable objective functions, making them unsuitable for the black-box optimization problems considered in this paper. Liu et al. [42] were the first to consider the coverage problem in the setting of black-box optimization. They propose a clustering-based swarm optimization algorithm (CluSO) for the coverage problem in the setting of multi-objective black-box optimization. In Section˜4, we provide empirical results demonstrating that our Bayesian optimization approach (MOCOBO) consistently outperforms CluSO by a large margin across all tasks considered.
3 Methods
@@
We consider the task of finding a set of $K<T$ solutions $S^{*}=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ that “covers" $T$ objectives $f_{1},\ldots,f_{T}$. Here, the set $S^{*}$ “covers" the $T$ objectives if, for each objective $f_{t}=f_{1},\ldots,f_{T}$, there is at least one solution $\mathbf{x}^{*}_{k}\in S^{*}$ for which $f_{t}$ is well optimized by $\mathbf{x}^{*}_{k}$. We evaluate how well a set of $K$ points $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}$ “covers" the $T$ objectives using the “coverage score" $c$:
$c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ $=\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (1)
@@
$S^{*}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq\mathcal{X}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ (2)
The coverage score in (1) only credits a single solution for each objective. If, for example, $f_{t}(\cdot)$ is maximized by one of the $\mathbf{x}_{i}$ in the solution set, improving the value of $f_{t}(\cdot)$ on some other $\mathbf{x}_{j\neq i}$ becomes irrelevant so long as $f_{t}(\mathbf{x}_{i})\geq f_{t}(\mathbf{x}_{j})$. In the setting where $K=1$, the coverage score collapses into a trivial linearization of the objectives and true multi-objective BO methods should be preferred. In the setting where $K=T$, the coverage score is trivially optimized by maximizing each objective independently.
@@ 3.1 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\left\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\right\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation˜2, we define $S^{*}_{D_{s}}$ as follows:
$S^{*}_{D_{s}}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}\sum_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k})$ (3)
$c(S^{*}_{D_{s}})$ denotes the best coverage score found by the optimizer after optimization step $s$.
@@ 3.1.1 Candidate selection with Expected Coverage Improvement (ECI)
Given $S^{*}_{D_{s}}$, our surrogate model’s predictive posterior $p(y\mid\mathbf{x},D)$ induces a posterior belief about the improvement in coverage score achievable by choosing to evaluate at $\mathbf{x}$ next. We extend the typical expected improvement (EI) acquisiton function in the natural way to the coverage optimization setting by defining expected coverage improvement (ECI):
$\mathrm{ECI}(\mathbf{x})$ $=\mathbb{E}_{p(y|\mathbf{x},D)}[\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})-c(S^{*}_{D_{s}}))]$ (4)
Here $c(S^{*}_{D_{s}})$ is the coverage score of the best possible covering set from among all data observed $D_{s}$ – as we shall see, constructing this set $S^{*}_{D_{s}}$ will be our primary challenge. $c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})$ is the coverage score of the best possible covering set after adding the observation $(\mathbf{x},\mathbf{y})$. Thus, ECI gives the expected improvement in the coverage score after making an observation at point $\mathbf{x}$. We aim to select points during acquisition that maximize ECI.
@@
$\mathrm{CI}(\mathbf{p}_{j})$ $=\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}}-c(S^{*}_{D_{s}})))$ (5)
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{p}_{j}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{j}$. We select and evaluate the candidate $\mathbf{p}_{j}$ with largest expected coverage improvement.
3.1.2 Greedy approximation of $S^{*}_{D_{s}}$
The candidate acquisition method in Section˜3.1.1 utilizes the best covering set of $K$ points among all data collected so far, $S^{*}_{D_{s}}$. On each step of optimization $s$, MOCOBO must therefore construct $S^{*}_{D_{s}}$ from all observed data $D_{s}$, as $S^{*}_{D_{s}}$ may change on each step of optimization after new data is evaluated and added to $D_{s}$.
Lemma 3.1 (NP-hardness of Optimal Covering Set).
Let $T,K$ be finite positive integers such that $K<T$. Let $f_{1},\ldots,f_{T}$ be real valued functions. Let $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ be a dataset of $n$ real valued data points such that for all $\mathbf{x}_{i}$, $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$. Let $S^{*}_{D_{s}}$ be the optimal covering set of size $K$ in $D_{s}$ as defined in Equation˜3. Then, constructing $S^{*}_{D_{s}}$ is NP-hard.
Proof.
See full proof of Lemma˜3.1 in Section˜A.9. ∎

### v4

@@ Abstract
In multi-objective black-box optimization, the goal is typically to find solutions that optimize a set of $T$ black-box objective functions, $f_{1},\ldots f_{T}$, simultaneously. Traditional approaches often seek a single Pareto-optimal set that balances trade-offs among all objectives. In contrast, we consider a problem setting that departs from this paradigm: finding a small set of $K<T$ solutions, that collectively “cover” the $T$ objectives. A set of solutions is defined as “covering” if, for each objective $f_{1},\ldots f_{T}$, there is at least one good solution. A motivating example for this problem setting occurs in drug design. For example, we may have $T$ pathogens and aim to identify a set of $K<T$ antibiotics such that at least one antibiotic can be used to treat each pathogen. This problem, known as coverage optimization, has yet to be tackled with the Bayesian optimization (BO) framework. To fill this void, we develop Multi-Objective Coverage Bayesian Optimization (MOCOBO), a BO algorithm for solving coverage optimization. Our approach is based on a new acquisition function reminiscent of expected improvement in the vanilla BO setup. We demonstrate the performance of our method on high-dimensional black-box optimization tasks, including applications in peptide and molecular design. Results show that the coverage of the $K<T$ solutions found by MOCOBO matches or nearly matches the coverage of $T$ solutions obtained by optimizing each objective individually. Furthermore, in in vitro experiments, the peptides found by MOCOBO exhibited high potency against drug-resistant pathogens, further demonstrating the potential of MOCOBO for drug discovery. All of our code is publicly available at the following link: https://github.com/nataliemaus/mocobo.
@@ 1 Introduction
Figure 1: Traditional multi-objective optimization for $T$ objectives might select any point along the Pareto frontier, but in some situations like this any Pareto optimal point performs poorly on at least one objective. In situations where multiple $K<T$ solutions are allowed ($\blacksquare$), we can sometimes optimize all objectives well. Note that this is a simplified schematic meant to illustrate intuition.
@@
Bayesian optimization (BO; [1, 2, 3]) is a general framework for sample-efficient optimization of black-box functions. By using a probabilistic surrogate model, such as a Gaussian process (GP; [4]), Bayesian optimization balances exploration and exploitation to identify high-performing solutions with a limited number of function evaluations. BO has been successfully applied in a wide range of domains, including hyperparameter tuning [5, 6], A/B testing [7], chemical engineering [8], drug discovery [9], and more.
For application domains such as drug discovery, however, the setup assumed by vanilla BO, where there is only one clear objective, is often insufficient. Consider the example of designing antibodies for treating a broad spectrum of pathogens. Here, the activity against each individual pathogen serves as an objective, each equally important. Ideally, one would seek to optimize all objectives simultaneously by obtaining a single broad-spectrum antibiotic. When this is not possible, however, we should seek antibiotics that are potent against a large enough group of pathogens. By identifying a set of such solutions—a covering set—we would obtain a set of drugs that are together lethal against the set of pathogens.
At first, this setup may be reminiscent of multi-objective optimization, which has been tackled under the black-box setup via multi-objective BO (MOBO; [10, 11, 12, 13, 14, 15, 16]). These methods search for Pareto-optimal solutions that balance the trade-offs between the individual objectives, enabling practitioners to post hoc select a single solution lying on the Pareto front based on their preferences. In cases where some objectives are completely incompatible with each other, however, selecting a single solution on the Pareto front cannot simultaneously satisfy performance requirements across those objectives. This is illustrated in Figure˜1, where no individual solution can provide acceptable performance across all objectives due to the extreme trade-offs between objectives.
Instead, our problem is distinctly described as the coverage optimization problem. More formally, consider a set of $T$ distinct objectives. Our goal is to obtain a set-valued solution of size $K<T$, where the set “covers” the $T$ objectives: For each objective, this $K$-element covering set should contain at least one solution that performs similarly to the maximizer of that objective. In the context of drug design, this translates to identifying $K$ drugs such that each of the $T$ pathogens is effectively addressed by at least one drug in the set. This setting is particularly relevant to applications where the size of $K$ is associated with downstream cost. For instance, in drug discovery, the number of compounds that can be synthesized is fundamentally constrained by the capacity of the facility.
Previously, the coverage optimization problem has been tackled in [17, 18, 19] through gradient-based algorithms, while for the black-box setup, Liu et al. [20] proposed an evolutionary algorithm. Over the years, variants of BO have demonstrated superior performance over evolutionary algorithms in high-dimensional [21, 22, 15] and multi-objective problems [14], which has been instrumental for its success in drug discovery [22, 23, 24, 25]. Therefore, it is natural to ask how we could harness the strengths and flexibility of BO to attack the coverage optimization problem.
In this work, we develop Multi-Objective Coverage Bayesian Optimization (MOCOBO)—a BO-based method for solving the coverage optimization problem with high-dimensional black-box objectives. MOCOBO primarily consists of a coverage-optimization analog of the celebrated expected improvement (EI; [26]) acquisition function, where the covering set is greedily constructed via submodular optimization [27]. We evaluate the resulting algorithm on realistic optimization tasks: drug discovery over molecules and peptides, and tuning of image processing pipelines. Baselines include the method of [20] and MOBO.
Our contributions are summarized as follows:
  1. $\bullet$
@@
    We propose MOCOBO, a coverage optimization algorithm for black-box objectives that can deal with structured, high-dimensional search spaces, and scale to large numbers of function evaluations.
  2. $\bullet$
@@ 4. $\bullet$
2 Background
Bayesian Optimization.
To solve the black-box optimization problem, BO operates by iteratively selecting a query according to a certain search policy [3] and observing the objective value on the query. The observations that have been obtained up to iteration $n\geq 0$ are collected into a dataset $D_{n}$ to form a surrogate model, typically GP, which supplies information to the policy. Most policies are chosen to maximize a function known as acquisition function [3] such that, at each step $n\geq 0$, the query $\mathbf{x}_{n+1}$ is found as
$\mathbf{x}_{n+1}=\operatorname*{arg\,max}_{\mathbf{x}\in\mathcal{X}}\alpha\left(\mathbf{x};D_{n}\right)\;,$ (1)
where EI [26] is a popular example of an acquisition function. In our case, the optimization problem is over covering sets. Naturally, an acquisition function for this setting is needed.
@@ 2 Background
For high-dimensional problems, the basic scheme of solving Equation˜1 is no longer effective. Local BO methods [21] address this issue by restricting the search space to a trust region that is adjusted adaptively. This simple modification has been shown to be effective and has been employed in various setups. In high-dimensional MOBO problems specifically, a variant of TuRBO [21] referred to as TuRBO-$M$ has been shown to be effective [14, 28]. TuRBO-$M$ works by running $M$ parallel local optimization runs, each maintaining its own dataset $D_{i}$ and surrogate model. Each local optimizer proposes candidates within a hyper-rectangular trust region $\mathcal{T}_{i}$, which is taken to be a rectangular subset of the space $\mathcal{X}$ centered on the incumbent $\mathbf{x}^{+}_{i}$. The side of each $\mathcal{T}_{i}$ is chosen to have a length $\ell_{i}\in[\ell_{\mathrm{min}},\ell_{\mathrm{max}}]$. If an optimizer improves the incumbent for $\rho_{\mathrm{succ}}$ consecutive iterations, $\ell_{i}$ expands to $\min(2\ell_{i},\ell_{\mathrm{max}})$. If not improved in $\rho_{\mathrm{fail}}$ iterations, $\ell_{i}$ is halved. Optimizers are restarted if $\ell_{i}$ drops below $\ell_{\mathrm{min}}$. We will later incorporate this scheme for solving the coverage optimization problem.
3 Multi-Objective Coverage Bayesian Optimization
3.1 Coverage Optimization
@@
Consider $T>1$ objectives $f_{1},\ldots,f_{T}$, each defined as $f_{t}:\mathcal{X}\rightarrow\mathbb{R}$ for each $t=1,\ldots,T$, all sharing the same input domain $\mathcal{X}$. For some user-defined parameter $K\in\{1,\ldots,T-1\}$, we consider the task of finding a set of $K$ solutions $S^{*}=\left\{\mathbf{x}^{*}_{1},\ldots,\mathbf{x}^{*}_{K}\right\}$ that “covers” the $T$ objectives ${\{f_{t}\}}_{t=1,\ldots,T}$. Specifically, the set $S^{*}\subseteq\mathcal{X}$ “covers” the $T$ objectives if, for each objective $f_{t}\in\{f_{1},\ldots,f_{T}\}$, there is at least one solution $\mathbf{x}^{*}_{k}\in S^{*}$ for which $f_{t}$ is well optimized by $\mathbf{x}^{*}_{k}$. We evaluate how well a set of $K$ points $\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}$ “covers" the $T$ objectives using the following “coverage score” Formally,
$c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})={\sum}_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k}).$ (2)
@@
$S^{*}$ $=\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq\mathcal{X}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}).$ (3)
The coverage score in (2) only credits a single solution for each objective. If, for example, $f_{t}$ is maximized by one of the $\mathbf{x}_{i}$ in the solution set, improving the value of $f_{t}$ on some other $\mathbf{x}_{j\neq i}$ becomes irrelevant so long as $f_{t}(\mathbf{x}_{i})\geq f_{t}(\mathbf{x}_{j})$. In the setting where $K=1$, the coverage score collapses into a trivial linearization of the objectives, and true multi-objective BO methods should be preferred. In the setting where $K=T$, the coverage score is trivially optimized by maximizing each objective independently. This problem has previously been formulated in [17, 18, 19, 20], where we will proceed by developing a new BO-based solution.
@@ 3.2 Multi-Objective Coverage Bayesian Optimization (MOCOBO)
In this section, we propose MOCOBO - an algorithm which extends Bayesian optimization to the problem setting above. On each step of optimization $s$, we use our current set of all data evaluated so far $D_{s}=\left\{(\mathbf{x}_{1},\mathbf{y}_{1}),\ldots,(\mathbf{x}_{n},\mathbf{y}_{n})\right\}$ to define the best covering set $S^{*}_{D_{s}}=\big\{\mathbf{x}^{*(s)}_{1},\ldots,\mathbf{x}^{*(s)}_{K}\in D_{s}\big\}$ found so far. Here $\mathbf{y}_{i}=(f_{1}(\mathbf{x}_{i}),\ldots,f_{T}(\mathbf{x}_{i}))$ and $n$ is the number of data points evaluated so far at step $s$. Following Equation˜3, we define $S^{*}_{D_{s}}$ as follows:
$S^{*}_{D_{s}}\quad=\quad\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}c(\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\})\quad=\quad\operatorname*{arg\,max}_{\left\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\right\}\subseteq D_{s}}{\sum}_{t=1}^{T}\max_{k=1}^{K}f_{t}(\mathbf{x}_{k}).$ (4)
Here, $c(S^{*}_{D_{s}})$ denotes the best coverage score found by the optimizer after optimization step $s$.
@@ 3.2.1 Candidate Selection with Expected Coverage Improvement (ECI)
Given $S^{*}_{D_{s}}$, our surrogate model’s predictive posterior $p(y\mid\mathbf{x},D)$ induces a posterior belief about the improvement in coverage score achievable by choosing to evaluate at $\mathbf{x}$ next. We naturally extend the typical expected improvement (EI; [26]) acquisition function to the coverage optimization setting by defining expected coverage improvement (ECI):
$\mathrm{ECI}(\mathbf{x})$ $=\mathbb{E}_{p(y|\mathbf{x},D)}[\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})-c(S^{*}_{D_{s}}))].$ (5)
Here $c(S^{*}_{D_{s}})$ is the coverage score of the best possible covering set from among all data observed $D_{s}$ – as we shall see, constructing this set $S^{*}_{D_{s}}$ will be our primary challenge. $c(S^{*}_{D_{s}\cup\left\{(\mathbf{x},\mathbf{y})\right\}})$ is the coverage score of the best possible covering set after adding the observation $(\mathbf{x},\mathbf{y})$. Thus, ECI gives the expected improvement in the coverage score after making an observation at point $\mathbf{x}$. We aim to select points during acquisition that maximize ECI. Note that the name ECI is shared with a method proposed by Malkomes et al. [29] for multi-objective experimental design, but their method targets an entirely different notion of “coverage.” (See Section˜5 for more details.)
@@
$\mathrm{CI}(\mathbf{p}_{j})$ $=\max(0,c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}}-c(S^{*}_{D_{s}}))).$ (6)
Here, $c(S^{*}_{D_{s}\cup\left\{(\mathbf{p}_{j},\hat{\mathbf{y}}_{j})\right\}})$ is the approximation of the coverage score of the new best covering set if we choose to evaluate candidate $\mathbf{p}_{j}$, assuming the candidate point will have the sampled objective values $\hat{\mathbf{y}}_{j}$. We select and evaluate the candidate $\mathbf{p}_{j}$ with the largest expected coverage improvement.
3.2.2 Greedy Approximation of Best Observed Covering Set